OpenStudy (anonymous):

The formula for the area of a trapezoid with altitude k and bases m and n is . A=k/2 (m+n) solve for m. I REALLYY need help!!!

OpenStudy (anonymous):

A=k/2 (m+n) 2A/k = m + n 2A/k - n = m m = 2A/k - n

OpenStudy (anonymous):

A) (2A-k) /k B) (2A+kn) / k C) (2A-n) / k

jimthompson5910 (jim_thompson5910):

\[\Large A = \frac{k}{2}(m+n)\] \[\Large 2A = k(m+n)\] \[\Large 2A = km+kn\] \[\Large 2A-kn = km\] \[\Large km = 2A-kn\] \[\Large m = \frac{2A-kn}{k}\]

OpenStudy (anonymous):

your a life saveerrr!!

jimthompson5910 (jim_thompson5910):

yw